# Noise-Aware Entanglement Generation Protocols for Superconducting Qubits with impedance-matched FBAR Transducers

### v2

@@ III.3 Entanglement Generation Results
$F_{TypeI}\sim\\ \frac{2(1-P_{th})^{2}}{(1-P_{th})^{2}(3-\eta_{l})+P_{th}(1-P_{th})(3-\eta_{l})+P_{th}^{2}(1-\eta_{l})}\\ \xrightarrow[P_{th}\rightarrow 0]{}\frac{2}{3-\eta_{l}}\\ F_{TypeII}\sim\frac{-(1-P_{th})^{2}}{P_{th}^{2}-2P_{th}-1}\xrightarrow[P_{th}\rightarrow 0]{}1\$ (73)
@@ IV Discussion and Outlook
Unlike OMC-based optomechanical transducers that exhibit single-photon optomechanical coupling rates of $10-100\;\mathrm{kHz}$ or more, FBAR transducers will have values of $<1\;\mathrm{kHz}$. While the lower coupling rate can be viewed as a disadvantage, our results show that it can be beneficial. The comparatively low optomechanical loading $R^{opt}_{EM}$ for the FBAR transducers at a given optical pump power requires larger matching capacitances and lower matching inductances, as compared to OMC-based transducers [27], which makes the design and fabrication of the matching networks more feasible. This allows the transducer’s piezo-mechanical components to have a higher static capacitance, enabling realistic fabrication targets, and allows for more accurate matching circuit designs factoring in non-idealities such as non-zero resistance $R_{L}$ and parasitic capacitances and inductances.
@@ VII Funding acknowledgments
Appendix A Adding transducer static resistance into Extended BVD model
@@
Appendix B $R_{L}\neq 0\;$from dielectric loss
@@
$R_{m}$ $44.588\;\mathrm{\Omega}$ $44.588\;\mathrm{\Omega}$
$\mathscr{L}_{+}^{2}$ $1$ $1$
$\mathscr{L}_{-}^{2}$ $41.303E-$ 4 $41.303E-$ 4
@@
$R_{OM,-}$ $5.959E-3\;\mathrm{\Omega}$ $5.959E-3\;\mathrm{\Omega}$
@@
$n_{o}$ $56.598E-$ 5 $56.598E-$ 5
$n_{th}$ $22.203E-$ 2 $22.203E-$ 2
Appendix C Matching inductor parasitic capacitance
The formalism developed the main text, like previous work, ignores stray capacitances associated with the matching inductor $L$ in the equivalent circuit model. This is despite the large matching inductances and small matching capacitances often required for OMC-based transducers [27], which may require large spiral inductors that have large stray, or parasitic, capacitances. The stray inductance of the matching capacitor should also be considered. In this section we examine the stray capacitances and inductances of the matching elements and their effect on impedance-matched transducer design.
Figure 12: A modified equivalent circuit for a transducer with parasitic terms incorporated into the matching network.
@@ Appendix C Matching inductor parasitic capacitance
Appendix D Bell-State Analysis for Time-Bin Photons
@@
Time bin qubits can be measured using a simple (balanced) Bell state analyzer with a non-polarizing 50:50 beamsplitter and two SNSPDs. With sufficient temporal resolution, this setup can be used to project a two-photon state onto |$\ket{\Psi^{-}}$ or |$\ket{\Psi^{+}}$. Here, the different Bell state projections correspond to different temporal detection patterns[74]. Specifically, if one detector registers a photon in the Early bin and the other detector registers a photon in the Late bin, the photons are projected into |$\ket{\Psi^{-}}$. The |$\ket{\Psi^{+}}$ state is projected when the same detector registers both Early and Late photons. The photonic Bell states, with early (E) and late (L) photons emerging from nodes A and B, take the form
@@
If we assume that the single photon detectors employed within the BSA can resolve the arrival time of photons with a given time-bin separation, we can proceed to use a balanced interferometer for Bell state measurements. Of course, this requires the dead time of the detector to be shorter than the time bin separation. For our chosen time-bin separation of $500\;\mathrm{ns}$, this requirement can be satisfied, as commercial SNSPDs have dead times falling in the $10-100\;\mathrm{ns}$ range. Assuming both |$\ket{\Psi^{-}}$ and |$\ket{\Psi^{+}}$ can be resolved, the BSA has a maximum efficiency of $50\%$.
Alternatively, we may use an unbalanced BSA, or a time-domain interferometer (TDI). One arm of the TDI has a delay line that applies a time delay equal to the time-bin separation, $\Delta t$, allowing for the erasure of which-time-bin information. Coincidence events occur when Early photons propagate through the long arm and Late photons propagate through the short arm. This will occur $50\%$ of the time for early-late input pairs. The other $50\%$ of the time, the Late photons will travel through the long arm, and vice versa, doubling the time delay between Early and Late. If we throw away non-coincident detection events, we are only able to distinguish the |$\ket{\Psi^{-}}$ state.
@@
Appendix E Blue-Detuned Transducer Operation
For completeness, here we consider entanglement heralding protocols that utilize spontaneous parametric down conversion (SPDC) in a blue-detuned transducer to generate entangled microwave-telecom photon pairs. This scenario is briefly explored using a formalism similar to out own in references [12, 59] but not in [46]. Blue-detuned operation with a Type-II protocol using time-bin photons, but not with a Type-I protocol, is explored in depth in reference [30].
For a Type-I protocol utilizing SPDC seeded by a blue-detuned transducer pump, each of two identical nodes produces entangled microwave-optical photon pairs. The microwave photon is converted to a superconducting qubit excitation via stimulated Raman absorption, while the optical photon is emitted by the transducer toward the BSA. The detection of an optical photon in the BSA (a single-click event) heralds the production of a single microwave photon. For a qubit initialized in its ground state, detection of an optical photon corresponds to the qubit being in its excited state. In a simplified picture, the qubit-cavity system begins in the state |$\ket{1_{m}}\ket{g}$, where $m$ denotes the microwave mode. Absorption of the photon via a Jaynes-Cummings interaction evolves the system to the state |$\ket{0_{m}}\ket{e}$.
@@ Appendix E Blue-Detuned Transducer Operation
|$\ket{\Psi}=\sqrt{P_{01}}\ket{g_{A}e_{B}}+\sqrt{P_{10}}\ket{e_{A}g_{B}}$ (87)
As discussed in [12, 29], this protocol is susceptible to additional source of readout infidelity. SPDC-based protocols allow for multi-photon generation events [59], where instead of a microwave optical pair |$\ket{1_{m}}\ket{1_{o}}$, the process generates |$\ket{N_{m}}\ket{N_{o}}$ pairs, where $N>1$. In the absence of number-resolving detectors, an N-photon generation event may satisfy the one-click condition of the protocol, generating a false positive event. Furthermore, the presence of multi-photon Fock states in the transducer’s microwave cavity may lead to multi-photon absorption events in the superconducting qubit, and other deleterious behaviors, such as qubit dephasing and spurious excitations due to residual population of the microwave cavity. Following [59], we list the infidelity contributions from multi-photon excitation events below.

### v3

@@ III.3 Entanglement Generation Results
$F_{TypeI}\sim\\ \frac{2(1-P_{th})^{2}}{(1-P_{th})^{2}(3-\eta_{l})+P_{th}(1-P_{th})(3-\eta_{l})+P_{th}^{2}(1-\eta_{l})}\\ \xrightarrow[P_{th}\rightarrow 0]{}\frac{2}{3-\eta_{l}}\\ F_{TypeII}\sim\frac{-(1-P_{th})^{2}}{P_{th}^{2}-2P_{th}-1}\xrightarrow[P_{th}\rightarrow 0]{}1\\$ (76)
@@ IV Discussion and Outlook
Unlike OMC-based optomechanical transducers that exhibit single-photon optomechanical coupling rates of $10-100\;\mathrm{kHz}$ or more, FBAR transducers will have values of $1\;\mathrm{kHz}$. While the lower coupling rate can be viewed as a disadvantage, our results show that it can be beneficial. The comparatively low optomechanical loading $R^{opt}_{EM}$ for the FBAR transducers at a given optical pump power requires larger matching capacitances and lower matching inductances, as compared to OMC-based transducers [27], which makes the design and fabrication of the matching networks more feasible. This allows the transducer’s piezo-mechanical components to have a higher static capacitance, enabling realistic fabrication targets, and allows for more accurate matching circuit designs factoring in non-idealities such as non-zero resistance $R_{L}$ and parasitic capacitances and inductances.
@@ VII Funding acknowledgments
Appendix A Comprehensive symbol table
Table 5: List of Symbols and Definitions
Symbol Units Definition
Mechanical Parameters
$\omega_{m}$ rad/s Mechanical mode frequency
$\omega_{s}$ rad/s Series resonant frequency
$\gamma_{m}$ rad/s Intrinsic mechanical linewidth
$C_{m}$ F Motional capacitance
$L_{m}$ H Motional inductance
$R_{m}$ $\Omega$ Motional resistance
$k_{eff}^{2}$ - Electromechanical coupling factor
$g_{EM}$ rad/s Electromechanical coupling rate
$\mathcal{C}_{EM}$ - Electromechanical cooperativity
Optical Parameters
$\omega_{o}$ rad/s Optical cavity frequency
$\kappa_{o}$ rad/s Total optical linewidth
$\kappa_{i}$ rad/s Intrinsic optical linewidth
$\kappa_{ext}$ rad/s External optical coupling rate
$g_{OM}$ rad/s Optomechanical coupling rate
$\mathcal{C}_{OM}$ - Optomechanical cooperativity
$\mathcal{C}_{OO}$ - Optical-optical cooperativity
$J$ rad/s Ring-ring coupling rate
$\eta_{o}$ - Optical external coupling efficiency
$\mathscr{L}_{\pm}^{2}$ - Lorentzian sideband amplitudes
$\Delta$ rad/s Optical pump detuning
Electrical/Matching Circuit
$C_{0}$ F Static capacitance
$R_{0}$ $\Omega$ Static resistance
$C_{T}$ F Matching capacitance
$L_{T}$ H Matching inductance
$R_{L}$ $\Omega$ Matching inductor resistance
$Z_{tx}$ $\Omega$ Transmission line impedance
$\omega_{LC}$ rad/s Matching circuit resonant frequency
$Q_{LC}$ - Matching circuit quality factor
$\kappa_{e}$ rad/s Electrical linewidth
$\eta_{e}$ - Electrical external coupling efficiency
Equivalent Circuit Elements
$R_{EM}$ $\Omega$ Electromechanical loading resistance
$R_{EM}^{opt}$ $\Omega$ Optimal electromechanical loading
$R_{OM}$ $\Omega$ Optomechanical coupling resistance
$R_{OM}^{eq}$ $\Omega$ Equivalent optomechanical resistance
$R_{OM,\pm}$ $\Omega$ Upper/lower sideband resistances
$R_{OO,\pm}$ $\Omega$ Optical-optical sideband resistances
Performance Metrics
$\eta$ - Conversion efficiency
$\eta_{alt}$ - Alternative efficiency expression
$n_{o}$ - Raman noise quanta per photon
$n_{th}$ - Thermal noise quanta per photon
$n_{m}$ - Thermal bath occupancy
$\Delta\omega$ rad/s Conversion bandwidth
Entanglement Protocol Parameters
$\Delta t$ s Transduction time duration
$r_{0}$ Hz Photon generation rate
$P_{1}$ - Single photon emission probability
$P_{th}$ - Thermal photon emission probability
$P_{01},P_{10}$ - Target Bell state probabilities
$F$ - Readout fidelity
$\tau_{ent}$ Hz Entanglement generation rate
$\eta_{l}$ - Link efficiency
$\eta_{det}$ - Detector efficiency
$t_{r}$ s Qubit reset time
Appendix B Adding transducer static resistance into Extended BVD model
@@
Appendix C $R_{L}\neq 0\;$rom dielectric loss
@@
$R_{m}$ $4.588\;\mathrm{\Omega} 44$$4.588\;\mathrm{\Omega}$
$\mathscr{L}_{+}^{2}$ 11 11
$\mathscr{L}_{-}^{2}$ $3\times 10^{-4} 1.303×10$$3\times 10^{-4}$
@@
$R_{OM,-}$ $59\times 10^{-3}\;\mathrm{\Omega} 5.959$$59\times 10^{-3}\;\mathrm{\Omega}$
@@
$n_{o}$ $6.598\times 10^{-5}$ $6.598\times 10^{-5}$
$n_{th}$ $203\times 10^{-2} 2.2$$203\times 10^{-2}$
Appendix D Matching inductor parasitic capacitance
In this section we examine the stray capacitances and inductances of the matching elements and their effect on impedance-matched transducer design.
@@ Appendix D Matching inductor parasitic capacitance
Appendix E Bell-State Analysis for Time-Bin Photons
@@
[…]
@@
Appendix F Building the Type-I conditional readout infidelity expression
In this section, we provide a detailed physical picture underlying the structure of the fidelity expressions given in Equations 66 and 67. An accompanying diagram for each scenario is shown in Fig. 13.
For the Type-I protocol, entanglement between remote nodes is established when one photon is detected at the Bell state analyzer. False positive events occur when one photon is detected, as desired, but the remote superconducting qubits are not in the desired Bell state. Let us review our notation:
Table 7: Notation for Type I Protocol Analysis
Sym Description Expression
$P_{1}$ Prob. a node emits a signal photon $1-e^{-r_{0}\Delta t}$
$P_{0}$ Prob. a node emits no signal photon $e^{-r_{0}\Delta t}$
$P_{\text{th}}$ Prob. a node emits a thermal photon $r_{0}n_{\text{th}}\Delta t$
$P_{01}$ Prob. Node A emits 1, Node B emits 0 $P_{1}P_{0}$
$P_{10}$ Prob. Node A emits 0, Node B emits 1 $P_{0}P_{1}$
$P_{00}$ Prob. both nodes emit 0 photons $P_{0}^{2}$
$\eta_{l}$ Link transmission prob.
Now we will break Equation 66 down term by term.
Term 1: $P1^{2}(1-P_{th})^{2}\eta_{l}(1-\eta_{l})$
Physical process: Both nodes emit signal photons (probability $P_{1}^{2}$), neither node emits a thermal photon (probability $(1-P_{th}^{2})$). One signal photon reaches a detector ($\eta_{l}$), and the other is lost before reaching a detector ($(1-\eta_{l})$).
Why it’s a false positive: The Bell State Analyzer sees one detector click, but both superconducting qubits are in their ground state, not in the desired $|eg\ket$$\pm\ket{eg}$ Bell state.
Term 2: $2P_{01}P_{th}(1-P_{th})\eta_{l}(1-\eta_{l})$
Physical process: One node emits a signal photon (probability $P_{01}$) and no thermal photon (probability $(1-P_{th})$), and the other node emits a thermal photon (probability $P_{th}$) but no signal photon. The thermal photon reaches a detector ($\eta_{l}$) but the signal photon is lost before reaching a detector ($(1-\eta_{l})$). A factor of two (assuming $P_{10}=P_{01}$) accounts for either node being the thermal emitter.
Why it’s a false positive: The Bell State Analyzer sees one detector click from the thermal photon but not the signal photon. From the detector click we make the correct inference that one signal photon was emitted, but the detected photon is uncorrelated with any superconducting qubit state. Without this correlation, the qubits remain in a separable product state rather than the desired Bell state. While the qubits are in orthogonal states, detecting a thermal photon fails to erase the ”which-path” information, which is needed to create a coherent Bell state between the remote qubits.
Term 3: $P_{00}P_{th}^{2}\eta_{l}(1-\eta_{l})$
Physical process: Neither node emits a signal photons (probability $P_{00}$). Both nodes emit thermal photons (probability $P_{th}$). One of the thermal photons reaches a photodetector ($\eta_{l}$), but the other is lost before reaching a detector ($(1-\eta_{l})$).
Why it’s a false positive: The Bell State Analyzer sees one detector click from thermal emission, but both superconducting qubits are in their excited state, not in the desired $|eg\ket$$\ket{eg}$ Bell state.
Term 4: $2P_{00}P_{th}(1-P_{th})\eta_{l}$
Physical process: Neither node emits a signal photon (probability $P_{00}$). One node emits a thermal photon (probability $P_{th}$), which reaches a photodetector ($\eta_{l}$). A factor of two accounts for either node being the thermal emitter.
Why it’s a false positive: The Bell State Analyzer sees one detector click from thermal emission, but both superconducting qubits are in their excited state, not in the desired $|eg\ket$$\ket{eg}$ Bell state.
Desired events: $2P_{01}(1-P_{th})^{2}\eta_{l}$
Physical process:
The overall conditional readout fidelity takes the form
$F=\frac{\text{(Desired events)}}{(\text{Desired events }+\text{ False positive events})}$ (90)
Figure 13: Diagrams of scenarios contributing (a-d) to Terms 1-4 of the infidelity expression and (e) desired entanglement heralding events.
Appendix G Blue-Detuned Transducer Operation
For completeness, here we consider entanglement heralding protocols that utilize spontaneous parametric down conversion (SPDC) in a blue-detuned transducer to generate entangled microwave-telecom photon pairs. This scenario is briefly explored using a formalism similar to our own in references [12, 62] but not in [47]. Blue-detuned operation with a Type-II protocol using time-bin photons, but not with a Type-I protocol, is explored in depth in reference [30].
[…]
@@ Appendix G Blue-Detuned Transducer Operation
[…]
@@
Appendix H Spatio-temporal properties of emitted optical modes
In this section, we take a focused look at the spatio-temporal properties of the photons emitted by the HBAR transducers, and the impact of fabrication imperfections on the achievable entanglement fidelity. Let us begin by considering the linewidths of the telecom photons that are emitted from the two transducers. An emitted photon’s linewidth (in frequency) is determined by the overall linewidth of the transducer’s optical cavity, $\Delta\nu$, which is inversely proportional to its quality factor $Q=\nu/\Delta\nu$. For the resonators studied in this work, the frequency separation between separate devices becomes comparable to the linewidth of the optical resonators themselves. Indeed, the linewidth of the micro-ring resonators ranges from about 3-10 pm, or 400 MHz to 1.2 GHz given a 193 THz frequency. This linewidth directly determines the temporal coherence properties of the emitted photons and the spatio-temporal mode structure of our entangled states.
Specifically, the optical cavity linewidth $\kappa\_o=\kappa\_i+\kappa\_{ext}\approx 150\;\mathrm{MHz}$ (from our Table 1 parameters) corresponds to a photon coherence time $\tau_{coh}\approx 1/(2\pi\kappa_{o})\approx 1\;\mathrm{ns}$. However, the actual duration of the emitted pulse is determined by the transduction process itself, which we found lasts about $\Delta t\approx 125\;\mathrm{ns}$, much longer than $\tau_{coh}$. This means our pulses are temporally long and spectrally narrow. Each emitted pulse is temporally extended over about 125 coherence times, but occupies only a small fraction ($1/\Delta t\approx 8\;\mathrm{MHz}$ out of $150\;\mathrm{MHz}$) of the available cavity bandwidth. The pulse is therefore single-mode in frequency, but temporally multi-mode in that it evolves over many cavity lifetimes.
For our entanglement heralding protocols, this has a few implications. First, for Type-I heralding, the entangled state is distributed across 125 temporal modes. The entanglement exists in a well-defined frequency mode, but is spread out over $125\;\mathrm{ns}$. Second, for Type-II heralding, we can still have confidence that we can distinguish between Early and Late time-bin photons, since the time-bin separation of $1.125\;\mathrm{\mu s}$ is much longer than both $\Delta t$ and $\tau_{coh}$.
[…]
Next we turn our attention to the spatial properties of the emitted photons. The pulse bandwidth corresponds to a coherence length of $\sim 37\;\mathrm{m}$, which is suitable for laboratory-scale quantum networking demonstrations. Active fiber phase stabilization will allow us to extend the range over which we can distribute entanglement using the photons emitted by these FBAR transducers.
We also include details about the spatial mode properties based on the transducer’s photonic circuit design, in case these are also desired. The Si3N4 ($n=2.00$ at 1550 nm) optical ring resonators in the FBAR devices have rectangular waveguides with typical sizes of 800 nm height $\times$ 2.1 $\mathrm{\mu m}$ width buried in an SiO2 ($n=1.44$ at 1550 nm) cladding. These devices support well-confined optical modes with a mode field diameter of approximately 1.6 $\mathrm{\mu m}$. Emitted signals are coupled to bus waveguides with a ring-bus gap on the order of $\sim 100\mathrm{{}^{\prime}s}\;\mathrm{nm}$ separation, then off-chip using edge couplers and attached high numerical aperture optical fibers.
Now let us explore the impacts of these parameters on the entanglement heralding protocols. Spatial mode mismatch between two network nodes may reduce interference visibility, impacting the fidelity ceiling of our protocols. We will define the normalized mode overlap as the integral
$Overlap=\left|\iint E_{1}^{*}(x,y)E_{2}(x,y)\,dx\,dy\right|^{2}\approx\left(\frac{2d_{1}d_{2}}{d_{1}^{2}+d_{2}^{2}}\right)^{2}$ (96)
with mode field diameters d1,d2d1,\;d2. For our Si3N4 ring resonator geometry, if we assume typical fabrication tolerances in our photonic waveguides, e.g. $\pm 10\;\mathrm{nm}$ width, $\pm 5\;\mathrm{nm}$ height, and $\pm 50\;\mathrm{nm}$ ring radius, we still anticipate a spatial mode variation of $\pm 0.5\%$. For small variations where $d1=d2+\Delta d$, this leads to an overlap of
$Overlap\approx 1-(\Delta d/d)^{2}\approx>99.999\%$ (97)
for a mode field diameter $d$ with variance $\Delta d$ between nominally identical transducer devices. In the frequency domain, if we assume a $\pm 5\%$ ($\pm 7.5\;\mathrm{MHz}$), this creates some spectral distinguishability. To maintain a $99\%$ fidelity ceiling, spatial mode mismatches of up to $\pm 2\%$ and spectral mismatches of $\pm 5\%$ are both acceptable. Combined with the narrow spectral bandwidth of the emitted photons, our approach should maintain high interference visibility for quantum local area networks.
Appendix I Optical photon distinguishability
One motivation behind generalizing the impedance matching formalism, as is done in the Main Text, is to generalize the design of impedance-matched devices to those including transducers with two optical modes, including double ring resonator devices. This is of paramount importance, because transducers with double ring resonators can have their hybridized optical modes frequency-tuned, e.g. via a DC bias voltage. This capability can be used to tune two transducers to emit photons with a high degree of spectral in-distinguishability. But a question remains: are single-ring FBAR transducers suitable for the proposed entanglement heralding protocols? For a few reasons, detailed below, they are.
Given the high-quality fabrication of Si3N4 integrated photonics available, optical micro-ring resonators for transducers, defined either on the same chip or on separate chips, have highly repeatable resonant features. As a specific example, an entanglement heralding experiment could be performed by separately connecting two remote superconducting qubit packages to two adjacent transducers on the same chip.We measured the optical resonances of many separate single-ring transducers on the same chip. See the table below as an example. Optical resonance values for six transducers, sampled arbitrarily throughout a chip with 56 devices, are listed.
Table 8: Device Resonances Closest to 1553 nm
Device Number Resonance Closest to 1553 nm
A 1553.0794 nm / 193.03099 THz
B 1553.3661 nm / 192.99537 THz
C 1553.4106 nm / 192.98984 THz
D 1553.3930 nm / 192.99202 THz
E 1553.1840 nm / 193.01799 THz
F 1553.3731 nm / 192.99450 THz
Of the devices listed here, Devices B and F have a resonant separation of 870 MHz, and Devices C and D have a resonant separation of 2.18 GHz. Devices A and E have a resonant frequency separation of 13 GHz. These frequency difference values fall well below the 100 GHz upper bound set by the jitter of our single photon detectors.
An additional consideration is the linewidths of the telecom photons that are emitted from the two transducers. The photon’s linewidth is determined by the overall linewidth of the transducer’s optical cavity, $\Delta\nu$, which is inversely proportional to its quality factor $Q=\nu/\Delta\nu$. For the resonators studied in this work, the frequency separation between separate devices becomes comparable to the linewidth of the optical resonators themselves. Indeed, the linewidth of the micro-ring resonators ranges from about 3-10 pm, or 400 MHz to 1.2 GHz given a 193 THz frequency. Two photons emitted from two of these transducers may therefore have non-negligible spectral overlap, rendering them effectively indistinguishable, irrespective of the discussion above.
But what if we can’t generate spectrally identical photons from two single-ring devices? Must we restrict ourselves to double-ring devices whose optical output can be frequency-tuned? As shown experimentally by Vittorini and Hucul et al. [79]), it is indeed possible to perform heralding gates between spectrally distinguishable flying optical photons. Their specific demonstration showed that if two quantum memories, A and B, emit photons with frequencies $\omega_{A}=\omega_{B}+\Delta\omega$, the remote memories can be projected into a Bell state using a standard Bell State analyzer. The frequency difference between the photons imposes a time-dependent phase on the resulting Bell state, which also depends on the time $\Delta t$ elapsed following Bell state generation:
[…]
The temporal resolution of the photon detectors determines the precision with which the relative phase can be tracked: if the temporal resolution $t_{r}$ satisfies $t_{r}\ll 2\pi/\Delta\omega$, the phase of the entangled memory state is well-defined. While it is preferred to have tunable optical resonators to achieve frequency-indistinguishable emission from two separate transducers, if specific conditions are met, heralding gates should succeed even with non-tunable single-ring transducers using a single pump laser. For superconducting nanowire single photon detectors with timing jitter values of about 10 ps, photons with frequency differences approaching 100 GHz can be accommodated within heralding gates. With proper phase tracking and quantum state tomography, our proposed protocols should be able to accommodate spectrally distinguishable photons within this limit.
In summary, with no mitigation, frequency mismatches exceeding the photon spectral bandwidth of $8\;\mathrm{MHz}$ will drop the quantum interference visibility to zero, dropping the fidelity to $50\%$, no better than classical correlations. With phase-tracked heralding, the fidelity ceiling for indistinguishable photons is preserved if the timing resolution is sufficient. As mentioned above, for an SNSPD timing jitter of $10\;\mathrm{ps}$, we have, at worst, a resolution requirement of $2\pi/\Delta\omega=77\;\mathrm{ps}$. This timing resolution is easily achievable for standard SNSPDs. Due to the experimental overhead associated with phase tracking, the timing jitter associated with the distinguishable photon pair may degrade the fidelity ceiling to $90\%$. Finally, with double-ring frequency tuning, spectral in-distinguishability should be maintained, so the anticipated fidelity ceiling should match the original baseline values of $90-98\%$.
Appendix J Sensitivity analysis of matching circuit errors
For realistic experiments, transducer devices, including impedance-matching superconducting circuits, will be subject to standard fabrication variances. Here, we analyze the impacts of fabrication variances of up to $\pm 20\%$ on the matching capacitor $C_{T}$ and matching inductor $L$, and calculated the impact of these variances on the transducer’s figures of merit, as well as the performance of the entanglement heralding protocol. Table 9 shows the variance in these figures of merit for worst-case $\pm 20\%$ variations in the matching network. Variances in transducer conversion efficiency are about $\sim 5\%$ for efficiency-maximized devices and $\sim 20\%$ for noise-minimized devices. Added thermal noise values are barely affected ($1\%$ change), and the conversion bandwidth decreases by about $3\%$ for efficiency-maximized devices and by about $16\%$ for noise-minimized devices.
Fortunately, the fidelity and entanglement rate of the Type-I and Type-II protocols is barely affected ($1\%$ change), even with noise-minimized transducers. A summary of the sensitivity analysis as compared to baseline values is included in Table 10.
Table 9: Sensitivity Analysis: ±20% Variations in Matching Components
Parameter Single Ring Double Ring
Max Eff Min Noise Max Eff Min Noise
Transducer Performance
Conversion Efficiency ($\Delta\eta$, %) 5.27-5.27 20.00-20.00 5.27-5.27 20.00-20.00
Thermal Noise ($\Delta n_{th}$, %) 0.03-0.03 0.61-0.61 0.03-0.03 0.61-0.61
Bandwidth ($\Delta BW$, %) 2.64-2.64 15.66-15.66 2.64-2.64 15.66-15.66
Type-I Protocol
Fidelity ($\Delta F_{I}$, %) 0.06-0.06 0.14-0.14 0.06-0.06 0.14-0.14
Rate ($\Delta R_{I}$, %) 0.59-0.59 0.51-0.51 0.59-0.59 0.51-0.51
Type-II Protocol
Fidelity ($\Delta F_{II}$, %) 0.59-0.59 0.35-0.35 0.59-0.59 0.35-0.35
Rate ($\Delta R_{II}$, %) 0.59-0.59 0.51-0.51 0.59-0.59 0.51-0.51
Table 10: Baseline Values and Sensitivity Analysis Summary
Parameter Single Ring Double Ring
Max Eff Min Noise Max Eff Min Noise
Baseline Values
Conversion Efficiency (%) $42.20$ $28.69$ $42.16$ $28.66$
Thermal Noise (quanta) $2.20\times 10^{-2}$ $6.11\times 10^{-3}$ $2.21\times 10^{-2}$ $6.11\times 10^{-3}$
Type-I Fidelity (%) $90.68$ $86.52$ $90.64$ $86.45$
Type-I Rate (kHz) $168.9$ $161.8$ $169.3$ $162.3$
Type-II Fidelity (%) $92.82$ $97.90$ $92.82$ $97.90$
Type-II Rate (kHz) $37.99$ $36.40$ $38.10$ $36.52$
Robustness Assessment
Efficiency Sensitivity Moderate High Moderate High
Noise Sensitivity Low Moderate Low Moderate
Protocol Fidelity Impact Minimal Minimal Minimal Minimal
Overall Robustness Good Fair Good Fair